# Bracketing Transduction Grammar BPE Byte Pair Encoding

### v1

@@
(eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version
11institutetext: CVSSP, University of Surrey
Guildford, United Kingdom
11email: {harry.walsh, b.saunders, r.bowden}@surrey.ac.uk
@@ Abstract
Sign Language Production (SLP) is a challenging task, given the limited resources available and the inherent diversity within sign data. As a result, previous works have suffered from the problem of regression to the mean, leading to under-articulated and incomprehensible signing. In this paper, we propose using dictionary examples and a learnt codebook of facial expressions to create expressive sign language sequences. However, simply concatenating signs and adding the face creates robotic and unnatural sequences. To address this we present a 7-step approach to effectively stitch sequences together. First, by normalizing each sign into a canonical pose, cropping, and stitching we create a continuous sequence. Then, by applying filtering in the frequency domain and resampling each sign, we create cohesive natural sequences that mimic the prosody found in the original data. We leverage a SignGAN model to map the output to a photo-realistic signer and present a complete Text-to-Sign (T2S) Sign Language Production (SLP) pipeline. Our evaluation demonstrates the effectiveness of the approach, showcasing state-of-the-art performance across all datasets. Finally, a user evaluation shows our approach outperforms the baseline model and is capable of producing realistic sign language sequences.
Keywords:
Sign Language Production Continuous Sequence Synthesis Human Pose Generation
@@ 1 Introduction
Analogous to the tone and rhythm used in spoken language, signed language exhibits prosody. The natural rhythm, stress and intonation that signed languages use to convey additional information [50]. Other phenomena, such as the impact of sentiment on the manual and non-manual features have also been studied, showing that emotion can change the form of a sign to exhibit distinctive profiles [35].
@@
Sign language corpora containing linguistic annotation are limited due to the cost and time required to create such annotations [23]. Previous works have attempted to directly regress a sequence of poses from the spoken language or representations such as gloss [36, 39, 14, 15, 52]. However, given that sign language is a low-resource language and the complexity is under-represented in small datasets, previous approaches have suffered from regression to the mean, resulting in under-articulated and incomprehensible signing. Additionally, previous works have implicitly modelled prosody, but due to the limited resources, it is often lost in production.
Figure 1: System overview of our approach to SLP. a) Spoken language to gloss, duration, facial token and cutoff Transformer. b) Stitcher that creates a pose sequence given a list of glosses, facial expressions, durations and a cutoff. c) The SignGAN module that produces a photo-realistic signer conditioned on a pose sequence.
@@
In this paper, we propose a novel approach to SLP that effectively stitches together dictionary examples to create a meaningful, continuous sequence of sign language. By using isolated signs, we ensure the sequence remains expressive, overcoming previous shortcomings related to regression to the mean. However, each example lacks non-manual features, so we propose a Noise Substitution Vector Quantization (NSVQ) transformer architecture to learn a dictionary of facial expressions that can be added to each sign to create a realistic sequence. To the best of our knowledge, we are the first to explicitly model aspects of signed prosody in the context of SLP. By training a translation model to predict glosses, alongside a duration, facial expression and a cutoff value, we can modify the sequence to eliminate robotic and unnatural movements. Resampling each sign to the predicted duration allows us to alter the velocity associated with signing stress and rhythm [49]. Furthermore, by applying filtering in the frequency domain, we can adjust the trajectory of each sign to create softer signing, akin to how signers modify a sign to convey sentiment [3, 13]. Our approach demonstrates it is capable of modifying the stitched sequence to emulate aspects of prosody seen in the original continuous data. Evaluation of the produced sequence with back translation showcases state-of-the-art performance on all three datasets.
@@
Furthermore, to conduct a realistic user evaluation we use SignGAN, a Generative Adversarial Network (GAN) capable of generating photo-realistic sign language videos from a sequence of poses [37]. Thus, we present a full Text-to-Sign (T2S) SLP pipeline that contains both manual and non-manual features. The user evaluation agrees that the approach outperforms the baseline method [38] and improves the realism of the signed sequence. An overview of the approach can be seen in Fig. 1.
Finally, when ground truth gloss timing is not available, such as on the RWTH-PHOENIX-Weather-2014T (PHOENIX14T) dataset, we leverage our stitching approach to perform sign segmentation. By creating an equivalent sequence and aligning it to the original data, we infer the duration of each sign.
The contributions of this paper can be summarised as:
  1. 1.
    A complete SLP pipeline, capable of generating expressive natural sign language videos from spoken language text.
  2. 2.
    A novel sign stitching approach to effectively join isolated examples and modify them to mimic sign prosody.
  3. 3.
    An approach to create a dictionary of non-manual facial features, that improves the realism of the produced sequence.
  4. 4.
    A novel approach to sign segmentation.
  5. 5.
    Comprehensive evaluation across several metrics and an in-depth user evaluation.
@@ 2.0.1 Sign language Translation:
For the last 30 years Computational Sign Language Translation (SLT) has been an active area of research [44]. Initially focusing on isolated Sign Language Recognition (SLR) where a single instance of a sign was classified using a Convolutional Neural Network (CNN) [24]. Subsequent works extended to Continuous Sign Language Recognition (CSLR), which requires both the segmentation of a video into its constituent signs and their respective classification [22]. Later Camgoz et al. introduced the task of Sign-to-Text (S2T) [5] using neural networks, an extension of CSLR that requires the additional task of translation to spoken language. S2T performance was later improved using a Transformer [47]. By using gloss as a supervision on the encoder with a Connectionist Temporal Classification (CTC) loss the model achieved state-of-the-art performance [6]. Although there has been a lot of work since, the architecture has since become the standard when computing back-translation performance [36, 38, 39].
2.0.2 Sign Language Production:
SLP is the reverse task to SLT, which aims to translate spoken sentences into continuous sign language. Early approaches to SLP used an animated avatar driven with either motion capture data or parameterised glosses [2, 8, 53, 9, 46, 10]. These works all required expensive annotation systems, such as the Hamburg Notation System (HamNoSys) [33] or SigML [18] and have shown to be unpopular with the Deaf community due to the robotic motion and under articulated signing [34]. None of these approaches attempt to join the isolated signs effectively. Instead opting to play each sign in sequence, with unnatural transitions in between. Furthermore, due to the effect called the ‘uncanny valley’, users do not feel comfortable with the resulting production [28]. Methods that use Motion Capture (MoCap) data achieve better results, but have a limited vocabulary due to the cost of collecting and annotating data [11]. Here, we directly extract skeletons from sign language videos, overcoming the need for expensive MoCap. While Mediapipe has become a standard for this task [26], it sometimes provides incorrect keypoint predictions, negatively impacting downstream tasks. To enhance the quality of pose estimation and uplift the prediction to 3D, Ivashechkin et al. proposed a method that combines Forward Kinematics with a neural network to ensure valid predictions [16]. In this work, we use this additional optimization to extract an accurate sign dictionary.
Research suggests that over 15 million sentence pairs are necessary for Neural Machine Translation (NMT) to outperform statistical approaches [21], but sign Language datasets with parallel text-gloss pairs are restricted to 50,000 sequences [23]. Consequently, sign languages can be classified as a low-resource language, making SLP a challenging task.
@@ 2.0.2 Sign Language Production:
Early deep learning SLP approaches used NMT and broke the task down into three steps, Text-to-Gloss (T2G), Gloss-to-Pose (G2P) and Pose-to-Sign (P2S) [43]. Saunders et al. introduced the Progressive Transformer (PT) [38], a transformer architecture that synthesises poses directly from text. Although better results were achieved using gloss as an intermediate representation, the approach suffered from regression to the mean, caused by the lack of training data and the diversity of lexical variants. To reduce the problem, adversarial training and Mixture Density Network (MDN) were applied [36, 39] and since then a range of approaches have been proposed [14, 15, 52]. However, visual inspection of the results shows that the approaches still suffer from regression to the mean, and as a result, they fail to effectively convey the translation. Here we propose a method to effectively join isolated signs, that does not suffer from regression to the mean. By filtering the movement in the frequency domain we ensure the produced sequence is stylistically cohesive. In addition, by modifying the length of each sign we recreate the prosody seen in the original data.
2.0.3 Vector Quantized Models:
The first Variational Autoencoders (VAE) model was introduced by Kingma et al. [20], but it struggled to capture fine-grain details. Later the first Vector Quantized Variational Autoencoders (VQ-VAE) architecture was introduced [30]. The model forced the latent space of the VAE to be discrete, and as a result, the model showed state-of-the-art image and audio generation performance. The model uses straight-though gradient estimation to propagate the gradient from the decoder to the encoder. This is necessary as the Vector Quantisation (VQ) operation uses an argmin operation to choose the closest matching codebook entry during training, and hence is non-differentiable. In total, the architecture required three losses to train. Kaiser et al. [17] reduced the required loss to two by adding an exponential moving average to update the codebook. Recently, Vali et al. [45] introduced the NSVQ VAE, this architecture further reduced the required loss functions to one. The technique approximated the quantization error and substituted it with normalised noise plus the original error. The model is trainable end-to-end, so we apply this approach to learn a dictionary of facial expressions.
2.0.4 Human Pose Synthesis:
Research indicates a strong preference among Deaf participants for sign language videos over skeleton representations [48]. Avatar approaches, due to their cartoon-like appearance, are also found to be unpopular, appearing more suitable for children rather than adults [51]. As a result, the field progressed to using photo-realistic approaches. Sanders et al. proposed the SignGAN architecture, a model capable of producing photo-realistic sign language videos, that are comprehensible by a native Deaf signer [37].
The model is designed as a GAN, trained with a combination of 5 losses. Once trained the generator network can synthesise pose and style conditioned images. We apply this model to conduct a user evaluation.
@@ 3 Methodology
SLP aims to facilitate the continuous translation from spoken to signed languages by converting a source spoken language sequence, $X=(x_{1},x_{2},...,x_{W})$ with W words into a video of photo-realistic sign, denoted as $V=(v_{1},v_{2},\ldots,v_{U})$ with U frames. To accomplish this we use two intermediate representations, following Fig. 1 from left to right. First, the spoken language is translated to a sequence of glosses, $Y=(y_{1},y_{2},\ldots,y_{G})$, face tokens, $\mathcal{F}^{t}=(f^{t}_{1},f^{t}_{2},\ldots,f^{t}_{G})$ and duration’s, $D=(d_{1},d_{2},\ldots,d_{G})$, all with length $G$. Additionally, for each sequence, we predict a low pass cutoff, $\mathcal{C}$ (Fig. 1.a). Each gloss and facial expression is stitched together using these parameters, to produce a continuous sequence of poses, denoted as $P=(p_{1},p_{2},...,p_{U})$ with U frames (Fig. 1.b). Finally, we use the pose sequence to condition the SignGAN module allowing us to produce a photo-realistic signer. Next, we provide a detailed explanation of each step in our pipeline, following the order illustrated in Fig. 1 from left to right. We then elaborate on the process of generating the cutoffs, gloss timestamps and the dictionary of facial expressions.
3.0.1 Translation Model
Given a spoken language sequence $X=(x_{1},x_{2},...,x_{W})$, our goal is to generate a corresponding sequence of glosses $Y=(y_{1},y_{2},...,y_{G})$. We design the transformer with four output layers, enabling the model to predict the corresponding duration (in frames) and facial expression for each gloss, plus a low-pass filter cutoff in Hz for each sequence. Thus the model learns the conditional probability $p(Y,D,\mathcal{F}^{t},C|X)$.
@@ 3.0.1 Translation Model
3.0.2 Stitching
For each dataset, we collect an isolated instance of each gloss in our target vocabulary. For each sign, we extract Mediapipe skeletons [26] and run an additional optimization to uplift the 2D skeletons to 3D [16]. The optimisation uses forward kinematics and a neural network to solve for joint angles, $J_{a}$. We choose to store our dictionary as joint angles, as this allows us to apply a canonical skeleton. This ensures the stitched sequence is consistent even if the original signers have different bone lengths. We define a dictionary of, $N_{s}$, signs as $DS=[S_{1},S_{2},...,S_{N_{s}}]$ where each sign in the dictionary consists of a sequence of angles, such that $S_{i}=(a_{1},a_{2},...,a_{U_{s}})$ and $a_{i}\in\mathbb{R}^{J_{a}}$, where $U_{s}$ is the duration in frames. In addition we define a learnt dictionary of, $N_{f}$, facial expressions as $DF=[F_{1},F_{2},...,F_{N_{f}}]$, where $F_{i}\in\mathbb{R}^{U_{f}\times J\times D}$. As illustrated in Fig. 3, the stitching pipeline is comprised of seven steps, we now detail each step.
@@ 3.0.2 Stitching
Thus, for each dictionary sign we create a sequence, $P^{\Delta}=(p^{\Delta}_{2},p^{\Delta}_{3},\ldots,p^{\Delta}_{U_{s}})$ $n\in{2,3,\ldots,U_{s}}$, representing the distance travelled for a dictionary sign. We remove the beginning frames once the sign has moved by a set threshold, $\alpha_{crop}$. The crop index is given by:
@@
To crop the end, we reverse the order of frames and repeat the process. However, for short sequences, this method may over-crop the sign, rendering it meaningless. We introduce an additional constraint $index_{end}-index_{start}<U_{min}$. If this condition holds, we reduce the threshold, $\alpha_{crop}$, by a factor of 0.1 and recalculate.
Step 4) As detailed above, we predict the duration of each gloss. Here, we utilize the duration to resample the length of each sign and facial expression, emulating the natural rhythm in the original data. This process involves upsampling or downsampling the sign using linear interpolation. However, due to co-articulation in the continuous data, some signs are predicted to be only a few frames long, which may not be sufficient for an isolated example. To address this, we impose a minimum sign length to ensure the produced sequence remains expressive. Once the facial expression and sign are resampled to the same length, we shift and rotate the face onto the body creating the complete skeleton.
@@
Step 6) We concatenate the signs and the stitched frames to form a single sequence. We then sum all the predicted durations and resample the sequence so it matches the ground truth.
Step 7) Finally, we apply a low-pass Butterworth filter to each keypoint over time [4]. The predicted cutoff value determines which frequencies are removed, and corresponds to the -3dB attenuation point. This step aims to enhance the stylistic cohesiveness of the sequence by smoothing out any sharp, quick movements not present in the original sequence. The transfer function can be formulated as;
$H(z)=\frac{1}{{1+\left(\frac{z}{{e^{j\cdot\omega_{c}}}}\right)^{2N}}}$ (5)
@@
3.0.3 SignGAN
Skeleton outputs have been shown to reduce Deaf comprehension compared to a photo-realistic signer [48]. Therefore, to gain valuable feedback from the Deaf community we train a SignGan model [37].
Given a pose sequence, $P=(p_{1},p_{2},\ldots,p_{U})$, generated by our stitching approach the model aims to generate the corresponding video of sign language, $V=(v_{1},v_{2},\ldots,v_{U})$ with $U$ frames. Each video frame is conditioned on the appearance of an individual extracted from a style image. For the user evaluation, we use the appearance of a single individual obtained by passing a style image through the style encoder.
3.0.4 Facial Expression Generation
To be truly understandable and accepted by the Deaf community non-manual features must be present in the final output. Here we are using a discrete sequence-to-sequence model to generate the translation. Therefore, we must learn a discrete vocabulary of facial expressions, $DF$, that can be added to the isolated signs. We apply a NSVQ to learn a spatial-temporal dictionary of facial expressions.
For each gloss in the original data, we extract the corresponding face mesh, denoted as $F$. We then resample each sequence to a constant length, $U_{f}$ and scale it to be a constant size. Signers in the dataset are often looking off center, therefore we normalize the average direction of the face so that it is looking directly forward. Similar to Fig. 2 (Encoder) we add positional encoding and then embed the sequence using a single linear layer. After the embedding is passed through the transformer encoder to the codebook. The NSVQ codebook learns a set of $N_{f}$ embeddings. We denote the embedded face sequence and therefore each codebook entry as $F^{z}_{i}\in\mathbb{R}^{U_{f}\times H}$, where $H$ is the embedding dimension. Each input is mapped to one codebook entry, the difference between the selected codebook entry and the input is then simulated using a normally distributed noise source. A product of the simulated noise and the encoder output is then passed to the decoder. We use the counter decoding technique from the PT [38], to drive the decoder. The decoder learns to reconstruct the original face sequence and the input counter values. Thus, the model is trained with the following loss function;

### v2

@@
[…]
University of Surrey
Guildford, UK Sign Stitching
@@ Abstract
Sign Language Production (SLP) is a challenging task, given the limited resources available and the inherent diversity within sign data. As a result, previous works have suffered from the problem of regression to the mean, leading to under-articulated and incomprehensible signing. In this paper, we propose using dictionary examples to create expressive sign language sequences. However, simply concatenating the signs would create robotic and unnatural sequences. Therefore, we present a 7-step approach to effectively stitch the signs together. First, by normalising each sign into a canonical pose, cropping and stitching we create a continuous sequence. Then by applying filtering in the frequency domain and resampling each sign we create cohesive natural sequences, that mimic the prosody found in the original data. We leverage the SignGAN model to map the output to a photo-realistic signer and present a complete Text-to-Sign (T2S) SLP pipeline. Our evaluation demonstrates the effectiveness of this approach, showcasing state-of-the-art performance across all datasets.
@@ 1 Introduction
Analogous to the tone and rhythm used in spoken language, signed language exhibits prosody. The natural rhythm, stress and intonation that signed languages use to convey information [Wilbur(2009)].
Figure 1: Overview of our approach. a) Spoken language to gloss, duration, facial expression and cutoff translation. b) Pose sequence generation. c) Photorealistic signer production.
@@
Sign language corpora containing linguistic annotation are limited due to the cost and time required to create such annotations [Konrad et al.(2020)Konrad, Hanke, Langer, Blanck, Bleicken, Hofmann, Jeziorski, König, König, Nishio, Regen, Salden, Wagner, Worseck, Böse, Jahn, and Schulder]. Previous works have attempted to directly regress a sequence of poses from the spoken language or representations such as gloss [Saunders et al.(2020a)Saunders, Camgoz, and Bowden, Saunders et al.(2021b)Saunders, Camgoz, and Bowden, Huang et al.(2021)Huang, Pan, Zhao, and Tian, Hwang et al.(2023)Hwang, Lee, and Park, Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu, Xie et al.(2024)Xie, Zhang, Taiying, Tang, Du, and Li, Stoll et al.(2020)Stoll, Camgoz, Hadfield, and Bowden, Tang et al.(2022)Tang, Hong, Guo, and Wang]. However, given that sign language is a low-resource language and the complexity is under-represented in small datasets, previous approaches have suffered from regression to the mean, resulting in under-articulated and incomprehensible signing. Additionally, previous works have implicitly modelled prosody, but due to the limited resources, it is often lost in production.
@@
In this paper, we propose a novel approach to SLP that effectively stitches together dictionary examples to create a meaningful, continuous sequence of sign language. By using isolated signs, we ensure the sequence remains expressive, overcoming previous shortcomings related to regression to the mean. However, each example lacks non-manual features, so we propose a Noise Substitution Vector Quantization (NSVQ) transformer architecture to learn a dictionary of facial expressions that can be added to each sign to create a realistic sequence. To the best of our knowledge, we are the first to explicitly model aspects of signed prosody in the context of SLP. By training a translation model to predict glosses, alongside a duration, facial expression and a cutoff value, we can modify the sequence to eliminate robotic and unnatural movements. Resampling each sign to the predicted duration allows us to alter the velocity associated with signing stress and rhythm [Wilbur(1999)]. Furthermore, by applying filtering in the frequency domain, we can adjust the trajectory of each sign to create softer signing, akin to how signers modify a sign to convey sentiment [Brentari(1998), Heloir and Gibet(2007), Reilly et al.(1992)Reilly, McIntire, and Seago]. Our approach demonstrates it is capable of modifying the stitched sequence to emulate aspects of prosody seen in the original continuous data. Evaluation of the produced sequence with back translation showcases state-of-the-art performance on all three datasets.
@@
Furthermore, to conduct a realistic user evaluation we use SignGAN, a Generative Adversarial Network (GAN) capable of generating photo-realistic sign language videos from a sequence of poses [Saunders et al.(2020b)Saunders, Camgoz, and Bowden]. Thus, we present a full Text-to-Sign (T2S) SLP pipeline that contains both manual and non-manual features. The user evaluation agrees that the approach outperforms the baseline method [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] and improves the realism of the signed sequence. An overview of the approach can be seen in Fig. 1.
@@ 2 Related Work
Sign language Translation: For the last 30 years Computational Sign Language Translation (SLT) has been an active area of research [Tamura and Kawasaki(1988)]. Initially focusing on isolated Sign Language Recognition (SLR) where a single instance of a sign was classified using a Convolutional Neural Network (CNN) [Lecun et al.(1998)Lecun, Bottou, Bengio, and Haffner]. Subsequent works extended to Continuous Sign Language Recognition (CSLR), which requires both the segmentation of a video into its constituent signs and their respective classification [Koller et al.(2015)Koller, Forster, and Ney]. Later Camgoz et al. introduced the task of Sign-to-Text (S2T) [Camgoz et al.(2018)Camgoz, Hadfield, Koller, Ney, and Bowden] using neural networks, an extension of CSLR that requires the additional task of translation to spoken language. S2T performance was later improved using a Transformer [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin].
Although there has been a lot of work since, the architecture has since become the standard when computing back-translation performance [Saunders et al.(2020a)Saunders, Camgoz, and Bowden, Saunders et al.(2020c)Saunders, Camgoz, and Bowden, Saunders et al.(2021b)Saunders, Camgoz, and Bowden].
Sign Language Production: SLP is the reverse task to SLT, which aims to translate spoken sentences into continuous sign language. Early approaches to SLP used an animated avatar driven with either motion capture data or parameterised glosses [Bangham et al.(2000)Bangham, Cox, Elliott, Glauert, Marshall, Rankov, and Wells, Cox et al.(2002)Cox, Lincoln, Tryggvason, Nakisa, Wells, Tutt, and Abbott, Zwitserlood et al.(2004)Zwitserlood, Verlinden, Ros, Van Der Schoot, and Netherlands, Efthimiou et al.(2012)Efthimiou, Fotinea, Hanke, Glauert, Bowden, Braffort, Collet, Maragos, and Lefebvre-Albaret, Van Wyk(2008), ElGhoul and Jemni(2011)]. These works all required expensive annotation systems, such as the Hamburg Notation System (HamNoSys) [Prillwitz and für Deutsche Gebärdensprache und Kommunikation Gehörloser(1989)] or SigML [Kennaway(2015)] and have shown to be unpopular with the Deaf community due to the robotic motion and under articulated signing [Rastgoo et al.(2021)Rastgoo, Kiani, Escalera, and Sabokrou]. None of these approaches attempt to join the isolated signs effectively. Instead opting to play each sign in the sequence, with unnatural transitions in between.
@@
Early deep learning SLP approaches used Neural Machine Translation (NMT) and broke the task down into three steps, Text-to-Gloss (T2G), Gloss-to-Pose (G2P) and Pose-to-Sign (P2S) [Stoll et al.(2020)Stoll, Camgoz, Hadfield, and Bowden]. Saunders et al. introduced the Progressive Transformer (PT) [Saunders et al.(2020c)Saunders, Camgoz, and Bowden], a transformer architecture that synthesises poses directly from text. Although better results were achieved using gloss as an intermediate representation, the approach suffered from regression to the mean, caused by the lack of training data and the diversity of lexical variants. To reduce the problem, adversarial training and Mixture Density Network (MDN) were applied [Saunders et al.(2020a)Saunders, Camgoz, and Bowden, Saunders et al.(2021b)Saunders, Camgoz, and Bowden] and since then a range of approaches have been proposed [Huang et al.(2021)Huang, Pan, Zhao, and Tian, Hwang et al.(2023)Hwang, Lee, and Park, Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu, Xie et al.(2024)Xie, Zhang, Taiying, Tang, Du, and Li, Tang et al.(2022)Tang, Hong, Guo, and Wang, Saunders et al.(2021a)Saunders, Camgoz, and Bowden]. However, visual inspection of the results shows that the approaches still suffer from regression to the mean, and as a result, they fail to effectively convey the translation. Here we propose a method to effectively join isolated signs, which means the produced sequences are guaranteed to be expressive and do not suffer from regression to the mean.
Furthermore, preliminary experiments reveal that each sign language sequence contains a distinct range of frequencies correlated to the signer’s style. Fast motions contain high frequencies, while soft, slow signing involves lower frequencies, typically within the range of 1 to 25 Hz. To emulate this characteristic we filter the produced sequences in the frequency domain using a low-pass Butterworth filter [Butterworth et al.(1930)]. This ensures the movements are stylistically cohesive. In addition, by adjusting the duration of each sign, we recreate the prosody observed in the original data.
@@ 3 Methodology
SLP aims to facilitate the continuous translation from spoken to signed languages by converting a source spoken language sequence, $X=(x_{1},x_{2},...,x_{W})$ with W words into a video of photo-realistic sign, denoted as $V=(v_{1},v_{2},\ldots,v_{U})$ with U frames. To accomplish this the approach uses two intermediate representations, following Fig. 1 from left to right. First, the spoken language is translated to a sequence of glosses, $Y=(y_{1},y_{2},\ldots,y_{G})$, face tokens, $\mathcal{F}^{t}=(f^{t}_{1},f^{t}_{2},\ldots,f^{t}_{G})$ and duration’s, $D=(d_{1},d_{2},\ldots,d_{G})$, all with length $G$. Additionally, for each sequence, we predict a low pass cutoff, $\mathcal{C}$, which we use to filter the movements (Fig. 1.a). Each gloss and facial expression is stitched together using these parameters, to produce a continuous sequence of poses, denoted as $P=(p_{1},p_{2},...,p_{U})$ with U frames (Fig. 1.b). Finally, we use the pose sequence to condition the SignGAN module allowing us to produce a photo-realistic signer. Next, we provide a detailed explanation of each step in our pipeline, following the order illustrated in Fig. 1 from left to right. We then elaborate on the process of generating the cutoff frequencies and the dictionary of facial expressions.
3.1 Translation Model
Given a spoken language sequence $X=(x_{1},x_{2},...,x_{W})$, our goal is to generate a corresponding sequence of glosses $Y=(y_{1},y_{2},...,y_{G})$. We design the transformer with four output layers, enabling the model to predict the corresponding duration (in frames) and facial expression for each gloss, plus a low-pass filter cutoff in Hz for each sequence. Thus the model learns the conditional probability $p(Y,D,\mathcal{F}^{t},\mathcal{C}|X)$.
@@ 3.1 Translation Model
3.2 Stitching
For each dataset, we collect an isolated instance of each gloss in our target vocabulary. For each sign, we extract Mediapipe skeletons [Lugaresi et al.(2019)Lugaresi, Tang, Nash, McClanahan, Uboweja, Hays, Zhang, Chang, Yong, Lee, et al.] and run an additional optimization to uplift the 2D skeletons to 3D [Ivashechkin et al.(2023)Ivashechkin, Mendez, and Bowden]. The optimisation uses forward kinematics and a neural network to solve for joint angles, $J_{a}$. We choose to store our dictionary as joint angles, as this allows us to apply a canonical skeleton. This ensures the stitched sequence is consistent even if the original signers have different bone lengths. We define a dictionary of, $N_{s}$, signs as $DS=[S_{1},S_{2},...,S_{N_{s}}]$ where each sign in the dictionary consists of a sequence of angles, such that $S_{i}=(a_{1},a_{2},...,a_{U_{s}})$ and $a_{i}\in\mathbb{R}^{J_{a}}$, where $U_{s}$ is the duration in frames. In addition we define a learnt dictionary of, $N_{f}$, facial expressions as $DF=[F_{1},F_{2},...,F_{N_{f}}]$, where $F_{i}\in\mathbb{R}^{U_{f}\times J\times D}$. Fig. 3 illustrates our seven-step stitching pipeline, we now detail each step.
@@ 3.2 Stitching
Thus, for each dictionary sign we create a sequence, $P^{\Delta}=(p^{\Delta}_{2},p^{\Delta}_{3},\ldots,p^{\Delta}_{U_{s}})$ $n\in{2,3,\ldots,U_{s}}$, representing the distance travelled for a dictionary sign. We remove the beginning frames once the sign has moved by a threshold, $\alpha_{crop}$. The crop index is given by:
@@
To crop the end, we reverse the order of frames and repeat the process.
Step 4) As detailed above, we predict the duration of each gloss. Here, we utilize the duration to resample the length of each sign and facial expression, emulating the natural rhythm in the original data. This process involves upsampling or downsampling the sign using linear interpolation. Once the facial expression and sign are resampled to the same length, we shift and rotate the face onto the body creating the complete skeleton.
@@
Step 6) We concatenate the signs and the stitched frames to form a single sequence. We then sum all the predicted durations and resample the sequence to match the ground truth.
Step 7) Finally, we apply a low-pass Butterworth filter to each keypoint over time [Butterworth et al.(1930)]. The predicted cutoff value determines which frequencies are removed, and corresponds to the -3dB attenuation point. This step aims to enhance the stylistic cohesiveness of the sequence by smoothing out any sharp, quick movements not present in the original sequence. The transfer function can be formulated as;
$H(z)=\left((1+\left(\frac{z}{{e^{j\cdot\omega_{c}}}}\right)^{2N}\right)^{-1}$ (5)
@@
3.3 SignGAN
Skeleton outputs have been shown to reduce Deaf comprehension compared to a photo-realistic signer [Ventura et al.(2020)Ventura, Duarte, and Giró-i Nieto]. Therefore, to gain valuable feedback from the Deaf community we train a SignGan model [Saunders et al.(2020b)Saunders, Camgoz, and Bowden]. Given a pose sequence, $P=(p_{1},p_{2},\ldots,p_{U})$, generated by our stitching approach the model aims to generate the corresponding video of sign language, $V=(v_{1},v_{2},\ldots,v_{U})$ with $U$ frames.
3.4 Facial Expression Generation
To be truly understandable and accepted by the Deaf community non-manual features must be present in the final output. Here we are using a discrete sequence-to-sequence model to generate the translation. Therefore, we must learn a discrete vocabulary of facial expressions, $DF$, that can be added to the isolated signs. We design a transformer base NSVQ architecture to learn a spatial-temporal dictionary of facial expressions. Using the ground truth gloss timing information, we extract the corresponding face mesh sequence, denoted as $F$. We then resample each sequence to a constant length, $U_{f}$ and scale it to be a constant size. Signers in the dataset are often looking off center, therefore we normalize the average direction of the face so that it is looking directly forward. Similar to Fig. 2 (Encoder) we add positional encoding and then embed the sequence using a single linear layer. After the embedding is passed through the transformer encoder to the codebook. The NSVQ codebook learns a set of $N_{f}$ embeddings [Vali and Bäckström(2022)]. We denote the embedded face sequence and therefore each codebook entry as $F^{z}_{i}\in\mathbb{R}^{U_{f}\times H}$, where $H$ is the embedding dimension. Each input is mapped to one codebook entry, the difference between the selected codebook entry and the input is then simulated using a normally distributed noise source. A product of the simulated noise and the encoder output is then passed to the decoder. We use the counter decoding technique from the PT [Saunders et al.(2020c)Saunders, Camgoz, and Bowden], to drive the decoder. The decoder learns to reconstruct the original face sequence and the input counter values. Thus, the model is trained end-to-end with the following loss function;